# Locate-then-edit for Multi-hop Factual Recall under Knowledge Editing

### v2

@@ Abstract
The locate-then-edit paradigm has shown significant promise for knowledge editing (KE) in Large Language Models (LLMs). While previous methods perform well on single-hop fact recall tasks, they consistently struggle with multi-hop factual recall tasks involving newly edited knowledge. In this paper, leveraging tools in mechanistic interpretability, we first identify that in multi-hop tasks, LLMs tend to retrieve implicit subject knowledge from deeper MLP layers, unlike single-hop tasks, which rely on earlier layers. This distinction explains the poor performance of current methods in multi-hop queries, as they primarily focus on editing shallow layers, leaving deeper layers unchanged. To address this, we propose IFMET, a novel locate-then-edit KE approach designed to edit both shallow and deep MLP layers. IFMET employs multi-hop editing prompts and supplementary sets to locate and modify knowledge across different reasoning stages. Experimental results demonstrate that IFMET significantly improves performance on multi-hop factual recall tasks, effectively overcoming the limitations of previous locate-then-edit methods.
@@ 1 Introduction
Large Language Models (LLMs) like ChatGPT (Achiam et al., 2023) and LLaMA-2 (Touvron et al., 2023) have emerged as powerful knowledge bases, demonstrating remarkable abilities in both factual knowledge representation and reasoning over complex queries (Etezadi & Shamsfard, 2022). However, as the need for updating and correcting knowledge within these models grows, research on knowledge editing (KE) has gained significant attention, focusing on cost-effective ways to modify specific information in LLMs (Mazzia et al., 2023). KE methods can be broadly classified into two categories based on whether they alter the original model weights: weight-preserving (Zhong et al., 2023) and weight-modifying approaches (Meng et al., 2022a; b). Weight-preserving methods aim to modify the model’s outputs by integrating external memory or leveraging strategies such as in-context learning without altering the underlying weights (Cheng et al., 2024b; a). In contrast, weight-modifying methods directly change the model’s internal weights to update the stored knowledge. Weight-modifying methods can be further categorized into learning-based and optimization-based methods. Learning-based methods update weights using gradients but face challenges such as overfitting and poor generalization. Optimization-based methods, such as ROME (Meng et al., 2022a) and MEMIT (Meng et al., 2022b), have introduced the “locate-then-edit” paradigm, which first identifies the knowledge storage layers and then adjusts their weights through optimization techniques to achieve the desired knowledge modification.
[…]
@@
Figure 1: (a) The existing locate-then-edit KE method updates new fact to the shallow layers of the model using a single-hop edit template. (b) For multi-hop fact recall tasks, especially when the edited fact is in the second or subsequent hops, the hops typically access the deeper layers which outputs the unmodified knowledge. (c) Our method introduces a prefix hop for each single-hop edit, creating a two-hop edit template. We utilize this new template to perform a furtherance edit, targeting the deeper layers for more effective knowledge updating.
@@
To answer this question, we first explored the mechanisms of the pre-edited model when handling single-hop and multi-hop factual recall tasks to gain insights. Using the example mentioned, we attempt to illustrate how the model reasons with the implicit subject “Spain” in $Q_{2}$, compared to the explicit mention in $Q_{1}$. We first use LogitLens (nostalgebraist, 2020; Dar et al., 2023) to interpret the information encoded in each layer’s hidden states by projecting them into the output vocabulary space. We find that at the last token position, the information of the implicit subject accumulates before the final answer, which is significantly different from the single-hop scenario. We then conduct causal intervention experiments (Li et al., 2024d) to further confirm the influence of the implicit subject on the final answer.
@@
Based on this, we further explore the mechanism of how the implicit subject influences the prediction of the final answer. By using causal intervention, our results indicate that in the multi-hop scenario, the implicit subject guides the emergence of the final answer by retrieving relevant knowledge from the later MLP layers. This contrasts sharply with the single-hop cases (Meng et al., 2022a; 2023), where the subject information is used to retrieve information from earlier MLP layers. Based on this difference, we provide an explanation for the unsatisfactory performance of the existing locate-then-edit methods for multi-hop tasks: Previous methods leveraging single-hop prompts for editing are insufficient as they only update the relevant knowledge in the shallow MLP layers but fail to propagate the changes to deeper layers. The model retains some of the old single-hop knowledge that only activated by additional implicit multi-hop fact recall mechanisms.
Based on these observations, we developed an advanced locate-then-edit KE method specifically designed to modify knowledge in both shallow and deep MLP layers, which we named Interpretability-Guided Furtherance Model Editing in a Transformer (IFMET). IFMET introduces a supplementary set for edit instances and generates multi-hop editing prompts, surpassing the limitations of single-hop prompts used in previous locate-then-edit approaches. This supplementary set helps us locate pre-existing knowledge that appears in later hops by leveraging the differences between the reasoning mechanism for single-hop and multi-hop queries. By leveraging each edit instance and its corresponding multi-hop editing prompt, IFMET locates and edits the knowledge stored in both earlier and later MLP layers, effectively addressing cases where the knowledge to be edited appears either in the first or subsequent hops during reasoning, as illustrated in Figure 1. Our contributions can be summarized as follows:
@@
  • Next, we pinpointed the components of the implicit subject that influenced the final answer within the later MLP layers. We demonstrated that the absence of edited knowledge of these components significantly impacted the model’s performance.
  • We propose IFMET, an advanced locate-then-edit KE method specifically designed to modify knowledge in both shallow and deep MLP layers using single and multi-hop edit prompts. Experimental results confirm the effectiveness of our method, showing that it successfully overcomes the limitations of previous locate-then-edit approaches in handling multi-hop factual recall tasks.
Due to the space limit, we refer readers to Appendix A for previous work.
@@ 2 Preliminaries
Notations. We define the set of knowledge as $\mathcal{K}=\{(s,r,o)\}\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}$, where $\mathcal{E}$ and $\mathcal{R}$ denote the set of entities and relations respectively. Each tuple $(s,r,o)\in\mathcal{K}$ represents that the corresponding entity of subject entity $s$ under relation $r$ is object entity $o$. An editing instance can be described in the form of a triplet: $e=\left(s,r,o\to o^{*}\right)$, where $o^{*}$ denotes the new edited object in place of the original object $o$ related to $s$ through $r$.
Multi-hop factual recall $Q$ requires multi-step reasoning to reach the final answer. Its reasoning process is composed of a chain of knowledge $C=\left(s_{1},r_{1},o_{1}\right)\oplus\dots\oplus\left(s_{n},r_{n},o_{n}\right)$, where $s_{1}$ is the start subject that is explicitly given in the question, $o_{n}$ is the final answer, and $\oplus$ used for chain adjacent reasoning steps which means the subject $s_{i+1}$ is identical to the object $o_{i}$ of preceding reasoning step. In order to better explore how the language model recalls multi-hop questions, we categorize the reasoning step into two types: explicit recall step $(s_{1},r_{1},o_{1})$ and implicit recall steps $\{\left(s_{2},r_{2},o_{2}\right),\dots,\left(s_{n},r_{n},o_{n}\right)\}$. The inference information required by the former subject $s_{1}$ explicitly appears in the prompt, while the subjects of the latter $s_{2}...s_{n}$ need to be inferred to obtain, which are called implicit subjects.
@@ 2.1 Factual Recall Tasks
Format of Factual Recall Tasks. Factual recall tasks refer to verifying whether the model $\mathcal{M}$ can correctly provide the final answer to a single-hop question or a multi-hop factual recall $Q$. Based on the two forms of declarative sentences and interrogative sentences, there are two different formats of factual recall tasks: Cloze-Format $Q_{cloze}$ and QA-Format $Q_{qa}$. For instance, given two-hop questions with the knowledge chain like (Paradiso, author, Dante Alighieri) $\oplus$ (Dante Alighieri, country of citizenship, Italy), $Q_{cloze}$ can be “The author of Paradiso is a citizen of”, while $Q_{qa}$ is “What country does the author of Paradiso hold citizenship in?”. If the model’s final answer is the same as the answer to the question, the recall is considered successful, which can be represented as $\mathcal{M}(Q_{cloze})=o_{n}$ or $\mathcal{M}(Q_{qa})=o_{n}$.
Multi-hop Factual Recall under Knowledge Editing. This task assesses whether the post-edited model can effectively leverage the updated knowledge for reasoning in multi-hop fact recall tasks. Given an edit $e=\left(s,r,o\to o^{*}\right)$, the edit prompt $T_{r}(s)$ and a chain of facts $C_{e}$ which includes $\left(s,r,o\right)$ as one of its components. After the post-edited model must leverage the new factual knowledge $\left(s,r,o^{*}\right)$ to answer the multi-hop query. For example, given edit (Paradiso, author, Dante Alighieri $\to$ Mark Twain), the model’s response of “The author of Paradiso is a citizen of” should change from the original answer Italy to the new answer USA.
@@ 2.2 Mechanistic Interpretation Tools
LogitLens. LogitLens (nostalgebraist, 2020) is a framework for interpreting the hidden states (activations) of language models such as GPT (Brown et al., 2020) by examining the logits (the raw prediction scores before they are transformed into probabilities) and corresponding probabilities. Specifically, for the hidden state $h^{i}_{l}$ at the $l$-th layer and position $i$, the logits $s^{i}_{l}$ and probabilities $p^{i}_{l}$ over the output vocabulary set $V$ are defined as follows:
@@
where $W_{U}$ denotes the unembedding matrix, which is the same matrix used in the final layer of the model for prediction. LogitLens aids in the decomposition of model predictions, elucidating the contributions from various input components such as MLPs and attention heads. This decomposition can be explored by modifying $h^{i}_{l}$ to the output from MLP $m^{i}_{l}$ or attention heads $a^{i}_{l}$, where $h^{i}_{l}=h^{l-1}_{i}+m^{i}_{l}+a^{i}_{l}$.  [^1] We employ GPT variants such as GPT-J Wang & Komatsuzaki (2021) that position attention in parallel to the MLP, which mathematically equates to models that calculate MLP sequentially after the attention module, as discussed in Brown et al. (2020). LogitLens posits that probabilities and logits provide insights into how the model prioritizes different potential tokens, as indicated by the proportion of related information. Specifically, we define $\text{Info}(h^{i}_{l},j)$ as the information related to token $j\in V$ contained in $h^{i}_{l}$, positively correlated with $s^{i}_{l}[j]$ and $p^{i}_{l}[j]$. To account for the probability variations across different layers, we define $\text{Info}(h^{i}_{l},j)$ as the layer-wise min-max normalized probability (Li et al., 2024d), where $L$ is the total number of layers:
@@
Causal Intervention on Hidden States. Causal intervention on hidden states Li et al. (2024d; a) involves deliberately altering specific hidden states in a model to observe the resulting changes in various metrics, thereby helping to establish cause-and-effect relationships. This process includes three pivotal components: the intervention operation $\mathcal{I}$ to be conducted, the target hidden state $\mathcal{H}$ selected for intervention, and the effect metric IE which measures the change caused by the intervention $\mathcal{I}$. In this paper, the possible hidden states $\mathcal{H}$ for intervention include the layer hidden states $h$, the output hidden states from MLPs $m$, and the output hidden states from attention heads $a$. We use the change in probability $p^{i}_{l}[j]$ from LogitLens as the effect metric IE, which quantifies the change in the probability of predicting the target token $j$ at layer $l$ for a specific position $i$. This metric enables us to determine whether specific components or tokens, have a causal influence on the model’s predictions.
@@
Figure 2: LogitLens results of the last token position at different layers. (a) Yellow line represents the information containing implicit subject $s_{2}$, i.e., $\text{Info}(h_{l},s_{2})$. Blue line represents the information for the final answer, i.e., $\text{Info}(h_{l},o_{2})$. (b) Yellow line represents the information of subject $s$. i.e., $\text{Info}(h_{l},s)$ and Blue line represents the information of the answer $o$, i.e., $\text{Info}(h_{l},o)$.
3 Mechanisms of Knowledge Storage and Reasoning
In this section, we will explore the reasoning mechanisms of the pre-edited model for both single-hop and multi-hop factual recall tasks. By comparing the knowledge utilization processes, we identify the reasons behind the suboptimal performance in multi-hop tasks and explain why the post-edited model tends to output the original answer instead of the new edited one. Specifically, we focus on two-hop tasks to better illustrate these distinctions. Experiments are conducted using a subset of single and two-hop data from MQuAKE-CF (Zhong et al., 2023) with the GPT-J (6B) model (Wang & Komatsuzaki, 2021). More detailed information about the data and experimental setup is provided in Appendix B.1.1.
3.1 How the Pre-edited Model Reasons Fact Recall Tasks

### v3

@@ Abstract
The locate-then-edit paradigm has shown significant promise for knowledge editing (KE) in Large Language Models (LLMs). While previous methods perform well on single-hop fact recall tasks, they consistently struggle with multi-hop factual recall tasks involving newly edited knowledge. In this paper, leveraging tools in mechanistic interpretability, we first identify that in multi-hop tasks, LLMs tend to retrieve knowledge with implicit subject information from deeper MLP layers, unlike single-hop tasks, which rely on shallow layers. This distinction explains the poor performance of current methods in multi-hop queries, as they primarily focus on editing shallow layers with single-hop edit prompts, leaving deeper layers unchanged. To address this, we propose IFMET, a novel locate-then-edit KE approach designed to edit both shallow and deep MLP layers. Beyond single-hop editing prompts, IFMET further incorporates multi-hop editing prompts to locate and modify knowledge across different stages of reasoning. Experimental results demonstrate that IFMET significantly improves performance on multi-hop factual recall tasks, overcoming the limitations of previous locate-then-edit methods.
Machine Learning, ICML
[…]
tcb@breakable
@@ 1 Introduction
Large Language Models (LLMs) like ChatGPT (Achiam et al., 2024) and LLaMA-2 (Touvron et al., 2023) have emerged as powerful knowledge bases, demonstrating remarkable abilities in both factual knowledge representation and reasoning over complex queries (Etezadi & Shamsfard, 2022). However, as the need for updating and correcting knowledge within these models grows, research on knowledge editing (KE) has gained significant attention, focusing on cost-effective ways to modify specific information in LLMs (Mazzia et al., 2023). KE methods can be broadly classified into two categories based on whether they alter the original model weights: weight-preserving (Zhong et al., 2023) and weight-modifying approaches (Meng et al., 2022a, b). Weight-preserving methods aim to modify the model’s outputs by integrating external memory or leveraging strategies such as in-context learning without altering the underlying weights (Cheng et al., 2024b, a). Weight-modifying methods can be further categorized into learning-based and optimization-based methods. The former update weights using gradients but face challenges such as overfitting and poor generalization. The latter, such as ROME (Meng et al., 2022a) and MEMIT (Meng et al., 2022b), have introduced the “locate-then-edit” paradigm, which first identifies the knowledge storage layers and then adjusts their weights through optimization techniques to achieve the desired knowledge modification.
Compared to weight-preserving methods and learning-based weight-modifying approaches, the locate-then-edit paradigm offers precise editing of the model’s internal knowledge with low computational costs (Zhang et al., 2024). However, despite the success of locate-then-edit methods in single-hop fact recall tasks (Li et al., 2024c), they share a common limitation (Zhong et al., 2023): The post-edited model struggles with multi-hop factual recall tasks involving the newly edited knowledge (see Table 3 for details). For example, after changing the knowledge(fact) “The capital of Spain” from “Madrid” to “Hartford”, the model correctly answers $Q_{1}=$ “What is the capital city of Spain?”. However, when posed with the multi-hop question $Q_{2}=$ “What is the capital city of the country where Pablo Picasso holds citizenship?”, it still responds with “Madrid” (Figure 1 (b)). This discrepancy raises a natural question: Has the locate-then-edit paradigm reached its limits for multi-hop factual recall tasks, or does it still hold unexplored potential?
@@
Figure 1: (a) The existing locate-then-edit KE method updates new fact to the shallow layers of the model using a single-hop edit prompt. (b) For multi-hop fact recall tasks, especially when the edited fact is in the second or subsequent hops, the hops typically access the deeper layers which outputs the unmodified knowledge. (c) Our method introduces a prefix hop for each single-hop edit, creating a two-hop edit prompt. We utilize this new prompt to perform a furtherance edit, targeting the deeper layers for more effective knowledge updating.
To address this question, we focus on the key distinction between the existing locate-then-edit paradigm and the multi-hop fact recall task, which lies in the single-hop editing prompts. Previous methods were primarily designed for single-hop factual recall tasks. They typically associate an edit instance, such as (Spain, Capital, Madrid $\to$ Hartford), with a single-hop editing prompt “The capital of Spain is” and use it for the following editing process. Considering this limitation, we propose the following hypotheses: 1) The recall process of the same fact may differ mechanistically between single-hop and multi-hop scenarios; 2) This discrepancy leads to the insufficiency of knowledge updates across the model, resulting in unsatisfactory performance on multi-hop factual recall tasks.
@@
To test hypothesis 1, we first explored the mechanisms of the pre-edited model when handling multi-hop and single-hop factual recall tasks. Using the example mentioned (Spain, Capital, Madrid), we attempt to illustrate how the model reasons with the implicit subject “Spain” in $Q_{2}$, compared to the explicit mention in $Q_{1}$. In Section 3.1, by interpreting the information encoded in each layer’s hidden states using LogitLens (nostalgebraist, 2020; Dar et al., 2023), We find that at the last token position, the information of the implicit subject accumulates before the final answer, which is significantly different from the single-hop scenario.
@@
We then investigate the causal influence of the implicit subject on the final answer and the mechanism by which it affects the prediction. By using causal intervention experiments (Li et al., 2024d), our results indicate that in the multi-hop scenario, the implicit subject causally guides the emergence of the final answer by retrieving relevant knowledge from the deeper MLP layers. This contrasts sharply with the single-hop cases (Meng et al., 2022a, 2023), where the subject information is used to retrieve information from shallow MLP layers. Based on this difference, We provide a more detailed explanation for hypothesis 2: Previous methods leveraging single-hop prompts for editing are insufficient as they only update the relevant knowledge in the shallow MLP layers but fail to propagate the changes to deeper layers. As a result, the deeper layers retain unedited knowledge that is only activated by implicit multi-hop fact recall mechanisms. Based on these observations, we developed an advanced locate-then-edit KE method specifically designed to modify knowledge in both shallow and deep MLP layers, which we named Interpretability-Guided Furtherance Model Editing in a Transformer (IFMET). To surpass the limitations of single-hop prompts, IFMET generates relevant multi-hop editing prompts for each edit instance. To address the issue of insufficient knowledge updates caused by differences in the reasoning mechanisms, IFMET extends existing methods by using multi-hop prompts for furtherance editing, effectively addressing cases in the single-hop and multi-hop scenario, as illustrated in Figure 1. Our contributions can be summarized as follows  [^1] Due to the space limit, we refer readers to Appendix A for previous work. :
@@
  • Next, we pinpointed the components of the implicit subject that influenced the final answer within the deeper MLP layers. We demonstrated that the absence of edited knowledge of these components significantly impacted the model’s performance.
  • We propose IFMET, an advanced locate-then-edit KE method specifically designed to modify knowledge in both shallow and deep MLP layers using single and multi-hop edit prompts. Experimental results confirm the effectiveness of our method, showing that it successfully overcomes the limitations of previous methods in handling multi-hop factual recall tasks.
@@ 2 Preliminaries
Notations. We define the set of knowledge(fact) as $\mathcal{K}=\{(s,r,o)\}\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}$, where $\mathcal{E}$ and $\mathcal{R}$ denote the set of entities and relations respectively. Each tuple $(s,r,o)\in\mathcal{K}$ represents that the corresponding entity of subject entity $s$ under relation $r$ is object entity $o$. An editing instance can be described in the form of a triplet: $e=\left(s,r,o\to o^{*}\right)$, where $o^{*}$ denotes the new edited object in place of the original object $o$ related to $s$ through $r$.
@@ 2.1 Factual Recall Tasks
Format of Factual Recall Tasks. Factual recall tasks refer to verifying whether the model $\mathcal{M}$ can correctly provide the final answer to a single-hop or multi-hop factual recall $Q$. $Q$ requires multi-step($\geq 1$) reasoning to reach the final answer. Its reasoning process is composed of a chain of knowledge $C=\left(s_{1},r_{1},o_{1}\right)\oplus\dots\oplus\left(s_{n},r_{n},o_{n}\right)$, where $s_{1}$ is the start subject that is explicitly given in the question, $o_{n}$ is the final answer. There are two different format question prompts for factual recall tasks: Cloze-Format $Q_{cloze}$ and QA-Format $Q_{qa}$. For instance, given two-hop questions with the knowledge chain like (Paradiso, author, Dante Alighieri) $\oplus$ (Dante Alighieri, country of citizenship, Italy), $Q_{cloze}$ can be “The author of Paradiso is a citizen of”, while $Q_{qa}$ is “What country does the author of Paradiso hold citizenship in?”. For better clarity, we categorize the multi-hop fact recall into two types: explicit recall step $(s_{1},r_{1},o_{1})$ and implicit recall steps $\{\left(s_{2},r_{2},o_{2}\right),\dots,\left(s_{n},r_{n},o_{n}\right)\}$. If the model’s final answer is the same as the answer to the question, the recall is considered successful, which can be represented as $\mathcal{M}(Q_{cloze})=o_{n}$ or $\mathcal{M}(Q_{qa})=o_{n}$.
Multi-hop Factual Recall under Knowledge Editing. This task assesses whether the post-edited model can effectively leverage the updated knowledge for reasoning in multi-hop fact recall tasks. Given an edit $e=\left(s,r,o\to o^{*}\right)$, the edit prompt $T_{e}$ and a chain of facts $C_{e}$ which includes $\left(s,r,o\right)$ as one of its components. The post-edited model must leverage the new factual knowledge $\left(s,r,o^{*}\right)$ to answer the multi-hop query. For example, given edit (Paradiso, author, Dante Alighieri $\to$ Mark Twain), the model’s response of “The author of Paradiso is a citizen of” should change from the original answer Italy to the new answer USA.
@@ 2.2 Mechanistic Interpretation Tools
LogitLens. LogitLens (nostalgebraist, 2020) is a framework for interpreting the hidden states (activations) of language models such as GPT (Brown et al., 2020). For the hidden state $h^{i}_{l}$ (token $i$ at the $l$-th layer), the logits $s^{i}_{l}$ and probabilities $p^{i}_{l}$ over the output vocabulary set $V$ are:
@@
where $W_{U}$ denotes the unembedding matrix used in the final prediction layer of the model. LogitLens also works for the decomposition of hidden states, such as MLPs $m^{i}_{l}$ and attention heads $a^{i}_{l}$, where $h^{i}_{l}=h^{i}_{l-1}+m^{i}_{l}+a^{i}_{l}$.  [^2] We employ GPT variants such as GPT-J (Wang & Komatsuzaki, 2021) that position attention in parallel to the MLP, which mathematically equates to models that calculate MLP sequentially after the attention module, as discussed in (Brown et al., 2020). LogitLens posits that probabilities and logits provide insights into how the model prioritizes different potential tokens, as indicated by the proportion of related information. So we define $\text{Info}(h^{i}_{l},j)$ as the information related to token $j\in V$ contained in $h^{i}_{l}$, positively correlated with $s^{i}_{l}[j]$ and $p^{i}_{l}[j]$. To account for the probability variations across different layers, we define $\text{Info}(h^{i}_{l},j)$ as the layer-wise min-max normalized probability (Li et al., 2024d), where $L$ is the total number of layers:
@@
Causal Intervention on Hidden States. Causal intervention on hidden states (Li et al., 2024d, a) involves deliberately altering specific hidden states in a model to observe the resulting changes in various metrics, thereby helping to establish cause-and-effect relationships. This process includes three pivotal components: the intervention operation $\mathcal{I}$ to be conducted, the target hidden state or its decomposition $\mathcal{H}$ selected for intervention, and the effect metric IE which measures the change caused by the intervention $\mathcal{I}$.
3 Mechanisms of Knowledge Storage and Reasoning
In this section, we will explore the reasoning mechanisms of the pre-edited model for multi-hop factual recall tasks and how they differ from those in single-hop. Specifically, we focus on two-hop tasks to better illustrate these distinctions, whose knowledge chain is represented as $C=\left(s_{1},r_{1},o_{1}\right)\oplus\left(s_{2},r_{2},o_{2}\right)$. In Section 3.1, we primarily investigate the differences in the recall mechanism of the same fact $(s_{2},r_{2},o_{2})$ when it severs as the implicit multi-hop step in the chain and the explicit single-hop reasoning step. In Sections 3.2, we explain why the model edited by existing method tends to output the original answer instead of the new edited one in multi-hop scenario [^3] All experiments in section 3 are conducted using a subset of single and two-hop data from MQuAKE-CF (Zhong et al., 2023) with the GPT-J (6B) model (Wang & Komatsuzaki, 2021). More detailed information about the data and the experimental setup is provided in Appendix B.2.1..
@@ 3 Mechanisms of Knowledge Storage and Reasoning
Figure 2: LogitLens results of the last token position at different layers. Yellow line represents the information containing implicit subject $s_{2}$, i.e., $\text{Info}(h_{l},s_{2})$. Blue line represents the information for the final answer, i.e., $\text{Info}(h_{l},o_{2})$. Larger versions of the sub-figures are available in Appendix Figure 8.
3.1 How Pre-edited Models Reason Fact Recall Tasks